# “We Demand Justice!”: Towards Social Context Grounding of Political Texts

### v2

@@ Abstract
Social media discourse frequently consists of ‘seemingly similar language used by opposing sides of the political spectrum’, often translating to starkly contrasting perspectives. E.g., “thoughts and prayers”, could express sympathy for mass-shooting victims, or criticize the lack of legislative action on the issue. This paper defines the context required to fully understand such ambiguous statements in a computational setting and ground them in real-world entities, actions, and attitudes. We propose two challenging datasets that require an understanding of the real-world context of the text. We benchmark these datasets against models built upon large pre-trained models, such as RoBERTa and GPT-3. Additionally, we develop and benchmark more structured models building upon existing Discourse Contextualization Framework and Political Actor Representation models. We analyze the datasets and the predictions to obtain further insights into the pragmatic language understanding challenges posed by the proposed social grounding tasks.
“We Demand Justice!”: Towards Social Context Grounding of Political Texts
Rajkumar Pujari Purdue University, USA rpujari@purdue.edu                        Chengfei Wu Purdue University, USA wu1491@purdue.edu                        Dan Goldwasser Purdue University, USA dgoldwas@purdue.edu
@@ 1 Introduction
Over the past decade micro-blogging websites have become the primary channel of engagement between politicians and people in the US, resulting in an increasingly succinct political discourse. Politicians from the same party often coordinate the phrasing of social messaging in developing events to amplify their impact (Vaes et al., 2011; Weber and Neumann, 2021). Hence, repetitive phrases such as ‘Thoughts and Prayers’ are extensively used to signal more nuanced stances across several events. The interaction among politicians from opposing parties often leads to messaging phrased similarly but intended to signal opposing real-world actions. ‘Thoughts and Prayers’ is used to express condolences by Republicans in mass shooting events, and by Democrats convey an angry or sarcastic tone as a call for action demanding ‘tighter gun control measures’. The example in fig. 1 shows contrasting interpretations of the phrase “We need to keep our teachers safe!” depending on different speakers and in the context of different events.
@@
Humans familiar with the stances of a politician and, possessing knowledge about the event from the news, can easily understand the intended meaning. But, computationally understanding the correct meaning of such language is challenging. Our main question in this paper is - Can an NLP model find the right meaning? From a linguistic perspective, we follow the distinction (Bach, 2008) between semantic interpretation (where the meaning is encoded in the utterance directly and does not change based on its external context), and pragmatic interpretation (that depends on extra-linguistic information). The latter has gathered significant interest in the NLP community recently (Bender and Koller, 2020; Bisk et al., 2020), focusing on language understanding, when grounded in an external context (Fried et al., 2023). To a large extent, the focus of such studies was on grounding language in a perceptual environment (e.g., image captioning (Andreas and Klein, 2016; Sharma et al., 2018; Alikhani et al., 2020), instruction following (Wang et al., 2016; Suhr et al., 2019; Lachmy et al., 2022), and game playing (Potts, 2012; Udagawa and Aizawa, 2019) tasks). Unlike these works, in this paper, we focus on grounding language in a social context, i.e., modeling the common ground (Clark and Brennan, 1991; Traum, 1994; Stalnaker, 2002) between the author and their social media followers that enables understanding what would otherwise be highly ambiguous utterances.
The Social Context Understanding, needed for building successful models for such tasks, can come from a wide variety of sources. The politician’s party affiliation, historical stances on the issue, social relationships, knowledge about the involved entities, and related prior and upcoming events, are a few examples of the relevant social context. In fig. 1 event #1#1\#1# 1, combining the event information (school shooting) with the speakers’ stances on gun control, would facilitate understanding the intended meaning of the text better.
@@
This paper operationalizes the problem of ‘Social Context Grounding’ by defining tasks capturing the intended meaning at different level of granularity. At the most basic level, providing the social context should help identify the entities targeted and the sentiment towards them. In fig. 1, the social context ⟨Ev.#1, Harris⟩and the text “we need to keep our teachers safe” $\Rightarrow$ negative attitude towards guns. A more nuanced account of meaning, which we formulate as a separate task, captures the specific means in which the negative attitude is expressed (the Interpretation in fig. 1). We present two datasets corresponding to these tasks, namely, ‘Target Entity and Sentiment Detection’ and ‘Vague Text Disambiguation’. In the first, the goal is to predict: $11$) whether a given entity is the intended target of a politician’s tweet and $22$) the sentiment towards the intended targets. We explicitly focus on tweets that do not always mention the targets in their text to incentivize modeling the pragmatic communicative intent of the text. In the second task, given an ambiguous political message such as “We demand justice” and its social context (associated event, & the author’s party affiliation), the task is to identify a plausible unambiguous explanation of the message. Note that the ground truth for all these tasks is based on human pragmatic interpretation, i.e., “guns” is a negative target of “we need to keep our teachers safe”, despite not being mentioned in the text, since it was perceived in this way by a team of human annotators reading the tweet and knowing social context. We show examples of each task in table 1. We describe the datasets in detail in section 3.
@@
We evaluate the performance of various models, as a way to test the need for social context and compare different approaches for modeling it. These include pre-trained LM-based classifiers, and LLM in-context learning (Brown et al., 2020a; Black et al., 2022), which use a textual representation of the social context. We adopt an existing graph-based discourse contextualization framework (Pujari and Goldwasser, 2021; Feng et al., 2022) to explicitly model the social context needed to solve the proposed tasks. Our results demonstrate that discourse contextualization models outperform other models on both tasks. We present an error analysis of the outputs to gain further insights. We describe the models in section 4 and the results in section 5.
@@
We also present qualitative visualization of a political event, Brett Kavanaugh Supreme Court Nomination in section 6.4 from the lens of target entity-sentiment to showcase a unique summary of the event discourse. Then, we perform human evaluation on ‘Vague Text Disambiguation’ and observe that humans find the task much easier than the evaluated models. We also present observations of human vs. LLM errors in disambiguations. Our contributions can be summarized as:
@@
    Evaluating various state-of-the-art context representation models on the task. Adopting existing discourse contextualization framework for the proposed tasks and evaluating GPT-3 in-context learning performance on the tasks
@@
    Performing human studies to benchmark the dataset difficulty and GPT-3 generation performance comparison with human workers [^1] Our data and code will be released under MIT license
@@ 3.1.1 Target-Sentiment Data Collection
We filter 3,45434543,4543 , 454 tweets for the three events using hashtags, keyword-based querying, and the dates of the event-based filtering from the Congress Tweets repository corpus [^2] https://github.com/alexlitel/congresstweets. We collect a subset of 1,77917791,7791 , 779 tweets that contain media (images/video) to increase the chances of the tweet text not containing the target entity mentions. Then, we use $66$ in-house human annotators and Amazon Mechanical Turk (AMT) workers who are familiar with the event context for annotation. We ask them to annotate the targeted entities and sentiments towards the targets. The authors of this paper also participated in the annotation process. We provide them with entity options based on the event in the focus of the tweet. Annotators are allowed to add additional options if needed. We also ask the annotators to mark non-targets for each tweet. We instruct them to keep the non-targets as relevant to the event as possible to create harder negative examples. Each tweet is annotated by three annotators. We filter $865865$ unique tweets with 5,89158915,8915 , 891 annotations, with majority agreement on each tweet. All the AMT annotations were additionally verified by in-house annotators for correctness. AMT workers were paid USD $11$ per tweet. It took $33$ minutes on average for each assignment, resulting in an hourly pay of USD $2020$. We include screenshots of the collection task GUIs in the appendix. We split the train, and test sets by events, authors, and targets to incentivize testing the general social grounding capabilities of the models. The test set also consists of authors, targets, and events not seen in the training set. We use Capitol Riots event for the test set of Target Entity and Sentiment Task. We split the examples into 4,37043704,3704 , 370 train, $511511$ development, and 1,00910091,0091 , 009 test examples.
@@ 3.2.1 Vague Text Data Collection
Data collection was done in several steps. (1)Vague Texts Collection. We collected vague text candidates from tweets by US politicians (i.e. senators and representatives) between the years $20192019$ to $20212021$ from Congress Tweets corpus. We identified a list of $99$ well-known events from that period and identified event-related tweets using their time frame and relevant hashtags. We used a pre-trained BERT-based (Devlin et al., 2019) NER model to collect tweets that do not contain any entity mentions to identify potential candidates for vague texts. We manually identified examples that could have contrasting senses by flipping their social context. We obtain $9393$ vague text candidates via this process.
(2)In-Context Plausible Meaning Annotation. We match the $9393$ ambiguous tweets with different events that fit them. We use both Democrat and Republican as the author party affiliation. We obtain $600600$ context-tweet pairs for AMT annotation. For each tweet, we ask AMT workers to annotate the following two aspects: 1) sentiment towards the three most relevant entities in the event (sanity check) and 2) a detailed explanation of the intended meaning given the event and author’s party affiliation. We obtain $469469$ reasonable annotations. After this step, each annotation was screened by in-house annotators. We ask three in-house annotators to vote on the correctness, appropriateness, and plausibility of the annotation given the context. Thus, we create a total of $374374$ examples.
(3)LLM-based Data Expansion. Using these examples, we further generate candidates for the task using LLM few-shot prompting. We use the examples from the previous step as in-context few-shot examples in the prompt. We use GPT-NeoX (Black et al., 2022) and GPT-3 (Brown et al., 2020a) for candidate generation. For each generated answer, manual inspection by three in-house annotators is performed to ensure data quality. We generate $928928$ candidates using GPT-NeoX and GPT-3. Human expert filtering results in $650650$ generations that pass the quality check. After removing redundant samples, we obtain $365365$ examples. Thus, we obtain a total of $739739$ annotations for this task. Then, for each of the $739739$ examples, we ask in-house annotators to select $33$ relevant negative options from the pool of explanations. We instruct them to pick hard examples that might contain similar entities as the correct interpretation. This results in 2,95629562,9562 , 956 binary classification data samples. We analyze and discuss the results of human validation of large LM generations in section 6.
@@ 4 Modeling Social Context
We evaluate GPT-3333gpt-3.5-turbo-1106 via OpenAI API in zero-shot and four-shot in-context learning paradigm on both tasks. We provide contextual information in the prompt in the form of short event descriptions and authors’ affiliation descriptions. Note that GPT-3 is trained on news data until Sep. $20212021$ which includes the events in our data (models denoted LLM Baseline).
@@ 4.1 Discourse Contextualization Framework
Discourse Contextualization Framework (DCF) (Pujari and Goldwasser, 2021) leverages relations among social context components to learn contextualized representations for text, politicians, events, and issues. It consists of encoder and composer modules that compute holistic representations of the context graph. The encoder creates an initial representation of nodes. Composer propagates the information within the graph to update node representations. They further define link prediction learning tasks over context graphs to train the model. They show that their representations significantly outperform several PLM-based baselines trained using the same learning tasks. Their qualitative evaluation shows that these representations capture interesting aspects of context information.
@@ 4.2 Political Actor Representation
Feng et al. (2022) propose the Political Actor Representation (PAR) framework, a graph-based approach to learn more effective politician embeddings. They propose three learning tasks, namely, $11$) Expert Knowledge Alignment $22$) Stance Consistency training & $33$) Echo chamber simulation, to infuse social context into the politician representations. They show that PAR representations outperform state-of-the-art models on Roll Call Vote Prediction and Political Perspective Detection.
@@ 5 Results
We show the results of our baseline experiments in tables 2 and 3. We evaluate our models using macro-averaged precision, recall, F1, and accuracy metrics. As the classes are not balanced, macro-F1 is the main metric we observe. Overall, several patterns emerge from the results consistent across all tasks. First, modeling social context is still an open problem. None of our models were able to perform close to human level. Second, adding context can help performance, compared to the No-Context baselines, models incorporating context performed better, with very few exceptions. Third, LLMs are not the panacea for social-context pragmatic tasks. Despite having access to a textual context representation as part of the prompt, and having access to relevant event-related documents during their training phase, these models under-perform compared to much simpler models that were fine-tuned for this task. Finally, explicit context modeling using the DCF model consistently leads to the best performance. Across all of our tasks, some variation of DCF-based model provided the best performance. The DCF model mainly represents the social context in the form of text documents for all nodes. Further symbolic addition of other types of context such as social relationships among politicians and relationships between various nodes could further help in achieving better performance on these tasks. In the Target Entity task, RoBERTa-base + DCF embeddings obtain $73.5673.56$ F1 vs. $68.8368.83$ for the best no-context baseline. Twitter bio and wiki-context hardly improve, demonstrating the effectiveness of modeling contextual information explicitly vs. concatenating context as text documents. No context performance well above the random performance of $5050$ F1 indicates the bias in the target entity distribution among classes. We discuss this in section 6.4. In Sentiment Identification task, we see that BERT-large + DCF back-propagation outperforms all the other models. Vague Text Disambiguation task results in table 3 show that DCF models outperform other models significantly. $71.7171.71$ F1 is obtained by BERT-base + DCF embeddings. BERT-base performing better than bigger PLMs might be due to DCF model’s learning tasks being trained using BERT-base embeddings.

### v3

@@ Abstract
Political discourse on social media often contains similar language with opposing intended meanings. For example, the phrase thoughts and prayers, is used to express sympathy for mass shooting victims, as well as satirically criticize the lack of legislative action on gun control. Understanding such discourse fully by reading only the text is difficult. However, knowledge of the social context information makes it easier.
We characterize the social context required to fully understand such ambiguous discourse, by grounding the text in real-world entities, actions, and attitudes. We propose two datasets that require an understanding of social context and benchmark them using large pre-trained language models and several novel structured models. We show that structured models, explicitly modeling social context, outperform larger models on both tasks, but still lag significantly behind human performance. Finally, we perform an extensive analysis, to obtain further insights into the language understanding challenges posed by our social grounding tasks.
“We Demand Justice!”:
Towards Social Context Grounding of Political Texts
Rajkumar Pujari  and Chengfei Wu  and Dan Goldwasser Purdue University, West Lafayette, USA {rpujari,wu1491,dgoldwas}@purdue.edu
@@ 1 Introduction
Over the past decade, micro-blogging websites have become the primary medium for US politicians to interact with general citizens and influence their stances for gaining support. As a result, politicians from the same party often coordinate the phrasing of their social messaging, to amplify their impact (Vaes et al., 2011; Weber and Neumann, 2021). Hence, repetitive, succinct phrases, such as “Thoughts and Prayers”, are extensively used, although they signal more nuanced stances. Moreover, the interaction among politicians from opposing parties often leads to messaging phrased similarly, but signaling opposing real-world actions. For example, ‘Thoughts and Prayers’, when used by Republicans, expresses condolences in mass shooting events, but when used by Democrats conveys an angry or sarcastic tone as a call for action demanding “tighter gun control measures”. Similarly, fig. 1 shows contrasting interpretations of the phrase “We need to keep our teachers safe!” depending on different speakers and in the context of different events.
@@
Humans familiar with the stances of a politician and, possessing knowledge about the event from the news, can easily understand the intended meaning of political phrases. However, automatically understanding such language is challenging. Our main question in this paper is - Can an NLP model find the right meaning? From a linguistic perspective, we follow the distinction (Bach, 2008) between semantic interpretation (i.e., meaning encoded directly in the utterance and does not change based on its external context), and pragmatic interpretation (that depends on extra-linguistic information). The latter has gathered significant interest in the NLP community recently (Bender and Koller, 2020; Bisk et al., 2020), focusing on language understanding, when grounded in an external context (Fried et al., 2023). To a large extent, the focus of such studies has been on grounding language in a perceptual environment (e.g., image captioning (Andreas and Klein, 2016; Sharma et al., 2018; Alikhani et al., 2020), instruction following (Wang et al., 2016; Suhr et al., 2019; Lachmy et al., 2022), and game playing (Potts, 2012; Udagawa and Aizawa, 2019) tasks). Unlike these works, in this paper, we focus on grounding language in a social context, i.e., modeling the common ground (Clark and Brennan, 1991; Traum, 1994; Stalnaker, 2002) between the author and their social media followers, that enables understanding an otherwise highly ambiguous utterances. The Social Context Understanding, needed for building successful models for such tasks, can come from a wide variety of sources. The politician’s affiliation and historical stances on the issue provide can capture crucial social context. Social relationships, knowledge about the involved entities, and related prior and upcoming events form important part of the puzzle as well. In fig. 1 event #1#1\#1# 1, combining the event information (school shooting) with the speakers’ gun control stances, would facilitate understanding the intended meaning of the text.
@@
The main motivation of this paper work is to operationalize the ‘Social Context Grounding’ problem as a pragmatic understanding task. From a practical perspective, this would enable the creation of better NLP-CSS models that can process social media text in settings that require contextualized understanding. We suggest several datasets, designed to evaluate this ability in computational models. These task capture the intended meaning at different level of granularity. At the most basic level, providing the social context can help identify the entities targeted, and the sentiment towards them. In fig. 1, the social context ⟨event#1, Harris⟩and the text “we need to keep our teachers safe” $\Rightarrow$ negative attitude towards guns. A more nuanced account of meaning, which we formulate as a separate task, captures the specific means in which the negative attitude is expressed (the Interpretation in fig. 1). We additionally present two datasets corresponding to these tasks, namely, ‘Target Entity and Sentiment Detection’ and ‘Vague Text Disambiguation’. In the first, the goal is to predict: $11$) whether a given entity is the intended target of a politician’s tweet and $22$) the sentiment towards the intended targets. We explicitly focus on tweets that do not always mention the targets in their text to incentivize modeling the pragmatic communicative intent of the text. In the second task, given an ambiguous political message such as “We demand justice” and its social context (associated event, & the author’s party affiliation), the task is to identify a plausible unambiguous explanation of the message. Note that the ground truth for all these tasks is based on human pragmatic interpretation, i.e., “guns” is a negative target of “we need to keep our teachers safe”, despite not being mentioned in the text, since it was perceived in this way by a team of human annotators reading the tweet and knowing social context. We show examples of each task in table 1. We describe the datasets in detail in section 3.
@@
We evaluate the performance of various models, as a way to test the need for social context and compare different approaches for modeling it. These include pre-trained LM-based classifiers, and LLM in-context learning (Brown et al., 2020a; Black et al., 2022), which use a textual representation of the social context. We also adopt an existing graph-based discourse contextualization framework (Pujari and Goldwasser, 2021; Feng et al., 2022), to explicitly model the social context needed to solve the proposed tasks. Our results demonstrate that the discourse contextualization models outperform other models on both tasks. We present an error analysis to gain further insights. We describe the models in section 4 and the results in section 5.
@@
We also present a qualitative visualization of a political event, Brett Kavanaugh Supreme Court Nomination (section 6.4), from target entity-sentiment perspective. It showcases a unique summary of the event discourse. We perform human evaluation on our ‘Vague Text Disambiguation’ dataset, and observe that humans find this task much easier than the evaluated models. We also present observations of human vs. LLM errors in disambiguation. In summary, our contributions are:
@@
    Evaluating various state-of-the-art context representation models on the task. We adopt existing discourse contextualization framework for the proposed tasks, and evaluate GPT-3’s in-context learning performance, as well.
@@
    Performing human studies to benchmark the dataset difficulty and GPT-3 generation performance, when compared to human workers. [^1] Our data and code is at https://github.com/pujari-rajkumar/language-in-context
@@ 3.1.1 Target-Sentiment Data Collection
We filter 3,45434543,4543 , 454 tweets for the three events using hashtags, keyword-based querying, and the dates of the event-based filtering from the Congress Tweets repository corpus [^2] https://github.com/alexlitel/congresstweets. We collect a subset of 1,77917791,7791 , 779 tweets that contain media (images/video) to increase the chances of the tweet text not containing the target entity mentions. Then, we use $66$ in-house human annotators and Amazon Mechanical Turk (AMT) workers who are familiar with the event context for annotation. We ask them to annotate the targeted entities and sentiments towards the targets. The authors of this paper also participated in the annotation process. We provide them with entity options based on the event in the focus of the tweet. Annotators are allowed to add additional options if needed. We also ask the annotators to mark non-targets for each tweet. We instruct them to keep the non-targets as relevant to the event as possible to create harder negative examples. Each tweet is annotated by three annotators. We filter $865865$ unique tweets with 5,89158915,8915 , 891 annotations, with majority agreement on each tweet. All the AMT annotations were additionally verified by in-house annotators for correctness. AMT workers were paid USD $11$ per tweet. It took $33$ minutes on average for each assignment, resulting in an hourly pay of USD $2020$. We include screenshots of the collection task GUIs in the appendix. We split the train, and test sets by events, authors, and targets to incentivize testing the general social grounding capabilities of the models. The test set also consists of authors, targets, and events not seen in the training set. We use Capitol Riots event for the test set of Target Entity and Sentiment Task. We split the examples into 4,37043704,3704 , 370 train, $511511$ development, and 1,00910091,0091 , 009 test examples. We compute the mean Cohen’s kappa score for annotations and report inter-annotator agreement for annotated targets (0.47) and sentiment (0.73)
@@ 3.2.1 Vague Text Data Collection
Data collection was done in several steps. (1)Vague Texts Collection. We collected vague text candidates from tweets by US politicians (i.e. senators and representatives) between the years $20192019$ to $20212021$ from Congress Tweets corpus. We identified a list of $99$ well-known events from that period and identified event-related tweets using their time frame and relevant hashtags. We used a pre-trained BERT-based (Devlin et al., 2019) NER model to collect tweets that contain few or no entity mentions to identify potential candidates for vague texts. We manually identified examples that could have contrasting senses by flipping their social context. We obtain $9393$ vague text candidates via this process.
(2) In-Context Plausible Meaning Annotation. We match the $9393$ ambiguous tweets with different events that fit them. We use both Democrat/Republican as the author party affiliation. We obtain $600600$ context-tweet pairs for AMT annotation. For each tweet, we ask AMT workers to annotate the following two aspects: 1) sentiment towards the three most relevant entities in the event (sanity check) and 2) a detailed explanation of the intended meaning given the event and author’s party affiliation. We obtain $469469$ reasonable annotations. After this step, each annotation was screened by in-house annotators. We ask three in-house annotators to vote on the correctness, appropriateness, and plausibility of the annotation given the context. Thus, we create a total of $374374$ examples.
(3)LLM-based Data Expansion. Using these examples, we further generate candidates for the task using LLM few-shot prompting. We use the examples from the previous step as in-context few-shot examples in the prompt. We use GPT-NeoX (Black et al., 2022) and GPT-3 (Brown et al., 2020a) for candidate generation. Manual inspection by three in-house annotators is performed for each generated answer to ensure data quality. We generate $928928$ candidates using GPT-NeoX and GPT-3. Manual filtering results in $650650$ generations that pass the quality check. After removing redundant samples, we obtain $365365$ additional examples. Thus, we obtain a total of $739739$ annotations for this task. Then, for each of the $739739$ examples, we ask in-house annotators to select $33$ relevant negative options from the pool of explanations. We instruct them to pick hard examples that potentially contain overlapping entities with the gold answer. This results in 2,95629562,9562 , 956 binary classification data samples. We analyze and discuss the results of human validation of large LM generations in section 6).
This process allows us to create three variants of the task: binary-classification, multiple-choice and generation variants. We evaluate several classification models on the binary classification variant (Tab.3). We evaluate LLMs on the generation variant (§6.2). We benchmark humans and the best models on the multiple-choice variant (§6.3).
@@ 4 Modeling Social Context
We evaluate GPT-3333gpt-3.5-turbo-1106 via OpenAI API in zero-shot and four-shot in-context learning paradigm on both tasks. We provide contextual information in the prompt as short event descriptions and authors’ affiliation descriptions. Note that GPT-3 is trained on news data until Sep. $20212021$ which includes the events in our data (models denoted LLM Baseline).
@@ 4.1 Discourse Contextualization Framework
Discourse Contextualization Framework (DCF) (Pujari and Goldwasser, 2021) leverages relations among social context components to learn contextualized representations for text, politicians, events, and issues. It consists of encoder and composer modules that compute holistic representations of the context graph. The encoder creates an initial representation of nodes. Composer propagates the information within the graph to update node representations. They define link prediction learning tasks over context graphs to train the model. They show that their representations significantly outperform several PLM-based baselines trained using the same learning tasks.
@@ 4.2 Political Actor Representation
Feng et al. (2022) propose the Political Actor Representation (PAR) framework, a graph-based approach to learn more effective politician embeddings. They propose three learning tasks, namely, $11$) Expert Knowledge Alignment $22$) Stance Consistency training & $33$) Echo chamber simulation, to infuse social context into the politician representations. They show that PAR representations outperform SOTA models on Roll Call Vote Prediction and Political Perspective Detection.
@@ 5 Results
The results of our baseline experiments are described in Tab. 2 and 3. We evaluate our models using macro-averaged precision, recall, F1, and accuracy metrics (due to class imbalance, we focus on macro-F1). Several patterns, consistent across all tasks, emerge. First, modeling social context is still an open problem. None of our models were able to perform close to human level. Second, adding context can help performance, compared to the No-Context baselines, models incorporating context performed better, with very few exceptions. Third, LLMs are not the panacea for social-context pragmatic tasks. Despite having access to a textual context representation as part of the prompt, and having access to relevant event-related documents during their training phase, these models under-perform compared to much simpler models that were fine-tuned for this task. Finally, explicit context modeling using the DCF model consistently leads to the best performance. The DCF model mainly represents the social context in the form of text documents for all nodes. Further symbolic addition of other types of context such as social relationships among politicians and relationships between various nodes could further help in achieving better performance on these tasks. In the Target Entity task, RoBERTa-base + DCF embeddings obtain $73.5673.56$ F1 vs. $68.8368.83$ for the best no-context baseline. Twitter bio and wiki-context hardly improve, demonstrating the effectiveness of modeling contextual information explicitly vs. concatenating context as text documents. No context performance well above the random performance of $5050$ F1 indicates the bias in the target entity distribution among classes. We discuss this in section 6.4. In Sentiment Identification task, we see that BERT-large + DCF back-propagation outperforms all other models. Vague Text Disambiguation task results in table 3 show that DCF models outperform other models significantly. $71.7171.71$ F1 is obtained by BERT-base + DCF embeddings. BERT-base performing better than bigger PLMs might be due to DCF model’s learning tasks being trained using BERT-base embeddings.